OpenStudy (anonymous):

What percent of 150 is 4.5?

OpenStudy (anonymous):

150-100 4.5 - x x = 4.5*100/150 = 450/150 = 45/15 = 3%

OpenStudy (anonymous):

or you could go 4.5/150 = 0.03 then multiply by 100 for your percent.